OpenStudy (anonymous):

how do you solve this problem? show me how to solve step by step. jack spends all of his allowance every week on candy. Last week he bought 3 candy bars and spent the remaining $2.75 on gum. This week he bought 5 candy bars and had $1.25 left for a big bag of skittles. how much allowance does jack get every week?

OpenStudy (anonymous):

Let C be the number of candy bars and A be Jack's allowance. last week's expenditure looks like A=3C+$2.75 this week A=5C+1.25

OpenStudy (anonymous):

This is an system of equations problem. last week's spending equals this weeks spending. or 3(price of candy bars) +2.75 = last week 5(price of candy bars) + 1.25 = this week last week = this week. What price of candy bars, which we will call x, satisfies both? 3x + 2.75 = allowance and 5x + 1.25 = allowance or 3x = allowance - 2.75 5x = allowance - 1.25 you should be able to take it from there...